A balloon that contains 0.500 L of helium at 25 °C is cooled to 11 °C, at a constant pressure. What volume does the balloon now occupy?
Answer #1

The answer is: 0.477 L. See solution below:

Ideal gas equation: PV = nRT

where P is pressure, V is volume, n is moles of gas, R is molar gas constant and T is temperature in K.

Since P, n and R are constant, V/T = constant

V1 = 0.500 L, V2 = ?

T1 = 25 °C = 298.15 K, T2 = 11 °C = 284.15 K

V1/T1 = V2/T2

0.500/298.15 = V2/284.15

Volume V2 = 0.477 L

answered by: Ms Cherie